OpenStudy (anonymous):

Which expression represents the composition [f o g o h](x) for the functions below? f(x) = –2x4 g(x) = 4x – 6 h(x) = x – 1

OpenStudy (anonymous):

I think it's -2(4x-10)^4, but I wanted to check.

OpenStudy (anonymous):

@Directrix

Directrix (directrix):

Let's work from the inside out. f(x) = –2x4 g(x) = 4x – 6 h(x) = x – 1 g of h(x) = g (x - 1) = 4* (x - 1) - 6 = 4x - 10 So, g(h(x)) = 4x - 10 and then ...

Directrix (directrix):

f(x) = –2x4 f(4x - 10) = -2 * (4x - 10)^4 = ? *That will be f(g(h(x))) = -2 * (4x - 10)^4 * I don't know if you are required to expand that.

OpenStudy (anonymous):

Nope, I'm not required to expand that. Thank you! I was wondering if you could clear up something for me? g(f(x))=[gof](x) Is this correct?

Directrix (directrix):

-2(4x-10)^4 is what you got which is what I got. This website is rocking and rolling like a ship in a tempest.

Directrix (directrix):

I don't use the circle notation myself because I have to think of the composition in a "layers of an onion" mindset. I would like to say that it is correct to write it that way but I cannot say so with certainty. Sorry.

OpenStudy (anonymous):

Ok, thank you. I've always used g(f(x)), but the algebra class I'm taking has been using the other way I showed you and I wasn't sure if it actually was another way to write it or if I was getting something mixed up.

Directrix (directrix):

The way the class is using the notation is easier on the eyes.

OpenStudy (anonymous):

Ah, I suppose. I'd better get on with my homework, considering I have a huge history test on Friday. Thank you for helping me! :)

Directrix (directrix):

[f o g o h](x) To keep the order of the composition straight in my head, I tend to use this nest: f(g(h(x)))

Directrix (directrix):

You are welcome.